OpenStudy (anonymous):

What is the first step in solving this equation? 4 + 2(x - 8) = 16 A. Add 4 to each side B. Divide each side by 2 C. Subtract 4 from each side D. Add 8 to both sides

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

What do you think?

OpenStudy (anonymous):

Obviously something to do with the parenthesis, but there is no answer on that there??

OpenStudy (anonymous):

I don't know.

OpenStudy (skullpatrol):

Why "obviously?"

OpenStudy (anonymous):

I have a poor choice of words. i am kinda angry at this question. XD

pooja195 (pooja195):

Look at the isolated number *hint hint* :P You don't always necessarily need to distribute first

OpenStudy (anonymous):

there is no god given way to solve this equation!

OpenStudy (anonymous):

you have some choices even if they think you do not

pooja195 (pooja195):

oh well we should ditch it :P

OpenStudy (anonymous):

Wait what?

OpenStudy (skullpatrol):

Get a clear idea in your mind what your GOAL is.

pooja195 (pooja195):

Look at the isolated number *hint hint* :P You don't always necessarily need to distribute first 4 + 2(x - 8) = 16 There is something else you can do without distributing

OpenStudy (anonymous):

Add 4+ 6?

OpenStudy (anonymous):

you should ditch the math teacher that wrote it you could distribute first for example you could also divide each side by 2, but that is not what they want you to say

pooja195 (pooja195):

Where did the 6 come from? .-.

pooja195 (pooja195):

You have the right idea but if we are trying to cancel the 4 what would we do?

OpenStudy (anonymous):

how would I ditch my math teacher? XD

pooja195 (pooja195):

You say good bye and simply leave :P

OpenStudy (anonymous):

and um it's B or C

OpenStudy (anonymous):

Well if the budget isn't in for my area. Then the school goes byebye. Maybe then I won't have to ditch him :P

pooja195 (pooja195):

Think: Why would you divide 2 and if you do what would you divide it by?

OpenStudy (anonymous):

He ditches me. R.i.P

pooja195 (pooja195):

im sure your parents will find a math teacher who can make harder problems : P

OpenStudy (anonymous):

XD. anyway So I would be dividing. but what am I diving exactly? is that the question? @pooja195

pooja195 (pooja195):

Nope

pooja195 (pooja195):

You dont need to divide by 2

OpenStudy (anonymous):

So how would I subtract then?

OpenStudy (anonymous):

Yes i am an idiot when it comes to math XD

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

WELL DONE CONGRATS :DDDD

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

I would subtract both sides by 4. so 8 - 4 = 4?

pooja195 (pooja195):

YES

pooja195 (pooja195):

wait no

OpenStudy (anonymous):

o

pooja195 (pooja195):

you subtract 4 from each SIDE so by 16

pooja195 (pooja195):



pooja195 (pooja195):

@AnimeWover do you understand?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

16 - 4 = 12?

OpenStudy (mathmale):

Solve 4 + 2(x - 8) = 16 for x.What would happen if you were to subtract 4 from both sides and then to simplify the result? Once you've done that, divide both sides by 2. Can you explain why this is needed? How would you then further isolate x? Compare these steps to the four given answer choices. Which will you choose?

OpenStudy (anonymous):

What do you mean by BOTH SIDES?

OpenStudy (anonymous):

nvm. i already got the answer. They are looking for the FIRST step. So the first step is to subtract 4 from both sides?

pooja195 (pooja195):

Yes